# Edward Witten

### v1

@@ 9.4 Further Developments
Once one identifies the entropy of a boundary region ${\sf A}$ as the generalized entropy of a quantum extremal surface of minimum generalized entropy, the possibility arises that a phase transition will occur in the entropy of a region if two quantum extremal surfaces exchange dominance. In our discussion of the Page curve, we encountered a phenomenon reminiscent of a phase transition: the entropy of an evaporating black hole is claimed to vary as a function of time in a way that becomes nonanalytic at the Page time as $G\to 0$ (see fig. 16). This has been interpreted as resulting from a phase transition in the quantum extremal surface P ; AEMM . Roughly speaking, prior to the Page time, the entropy of an evaporating black hole is described in terms of an empty quantum extremal surface, and after the Page time it is described in terms of a quantum extremal surface located near the black hole horizon. This is perhaps the most significant recent result in the subject, but explaining it here would take us too far afield. See malda ; malda2 for expositions.
@@ 10 What is a White Hole?
Figure 30: In (a), a star collapses to form a black hole. The diagonal black line labeled $H$ is the future horizon of an outside observer. The wiggly line at the top represents the black hole singularity. ${\rm i}^{-}$ and ${\rm i}^{+}$ are past and future timelike infinity for an observer outside the horizon. Sketched in (b) is the time-reversed spacetime: a white hole spontaneously emits a star. The diagonal black line labeled $H$ is the past horizon of an outside observer – the region that the observer cannot enter or influence. ${\rm i}^{-}$ and ${\rm i}^{+}$ are past and future timelike infinity for an observer outside the horizon. The diagonal line $\gamma$ is the worldline of an infalling massless particle that was injected from a great distance far in the past. Initially, its energy was of of order the Hawking temperature; by the time it collides with the star that is emerging from the white hole, its energy is exponentially large. This exponential blueshift is the time-reversal of the exponential redshift on which Hawking’s derivation of black hole evaporation is based.
@@
Acknowledgements I thank Z. Komargodsky, J. Maldacena, and D. Stanford for discussions, and A. Herdershee, T. Jacobson, B. Zhao, and especially R. Wald for comments on the manuscript. Research supported in part by NSF Grant PHY-2207584.

### v4

@@ 9.4 Further Developments
Once one identifies the entropy of a boundary region ${\sf A}$ as the generalized entropy of a quantum extremal surface of minimum generalized entropy, the possibility arises that a phase transition will occur in the entropy of a region if two quantum extremal surfaces exchange dominance. In our discussion of the Page curve, we encountered a phenomenon reminiscent of a phase transition: the entropy of an evaporating black hole is claimed to vary as a function of time in a way that becomes nonanalytic at the Page time as $G\to 0$ (see fig. 16). This has been interpreted as resulting from a phase transition in the quantum extremal surface P ; AEMM . Roughly speaking, prior to the Page time, the entropy of an evaporating black hole is described in terms of an empty quantum extremal surface, and after the Page time it is described in terms of a quantum extremal surface located near the black hole horizon. This is perhaps the most significant recent result about black hole thermodynamics, but explaining it here would take us too far afield. See malda ; malda2 for expositions.
@@ 10 What is a White Hole?
Figure 30: In (a), a star collapses to form a black hole. The diagonal black line labeled $H$ is the future horizon of an outside observer. The wiggly line at the top represents the black hole singularity. ${\rm i}^{-}$ and ${\rm i}^{+}$ are past and future timelike infinity for an observer outside the horizon. Sketched in (b) is the time-reversed spacetime: a white hole spontaneously emits a star. The diagonal black line labeled $H$ is the past horizon of an outside observer – bounding the region that the observer cannot enter or influence. ${\rm i}^{-}$ and ${\rm i}^{+}$ are past and future timelike infinity for an observer outside the horizon. The diagonal line $\gamma$ is the worldline of an infalling massless particle that was injected from a great distance far in the past. Initially, its energy was of of order the Hawking temperature; by the time it collides with the star that is emerging from the white hole, its energy is exponentially large. This exponential blueshift is the time-reversal of the exponential redshift on which Hawking’s derivation of black hole evaporation is based.
@@
Acknowledgements I thank Z. Komargodsky, J. Maldacena, and D. Stanford for discussions, and A. Herdershee, T. Jacobson, B. Zhao, and especially R. Wald for comments on the manuscript. Research supported in part by NSF Grant PHY-2207584. No data were generated in the course of this work.